# Knowledge Prompting: How Knowledge Engineers Use Large Language Models

### v1

@@ 1. Introduction
    What factors may affect knowledge engineers’ trust and uptake of the LLM technology?
Our contributions are threefold:
  • we present one of the first user studies on the interaction of knowledge engineers with LLMs
  • we identify key areas of strength and challenge for KE with LLMs
  • we propose data cards for KGs.
Our findings:
  • give direction to designers of KE AI copilots to create responsible applications
  • guide KE researchers and practitioners developing new methodologies
  • advise practitioners wishing to use advanced AI technologies responsibly.
2. Background and Related Work
In this section, we describe the background and related work for KE, automation in KE, and responsible AI.
2.1. Knowledge engineering
KGs represent real-world entities and their relations (Hogan et al., 2022). The graph structure offers efficient data management, availability of data in different modalities, use of machine learning and crowdsourcing techniques, and lowers cost. KGs can be built from numerous sources and often use a schema, an ontology or rules for representation (Hogan et al., 2022).
KG construction activities workflow depends on the actors involved, the represented domain, the construction tools, and the available data source (Hogan et al., 2022). Key actors generally include domain experts, specialising in the intricacies of the topic under consideration. Knowledge engineers elicit insights from these experts to create a conceptual domain model, and in turn, ontology engineers represent this knowledge in a suitable machine language (Simperl and Luczak-Rösch, 2014). KGs can be built and enriched manually using human collaboration (e.g., enterprise employees or crowdsourcing platforms) (Vrandečić and Krötzsch, 2014; He et al., [n. d.]), and automatically (e.g., NLP) (Maynard et al., 2017; Martinez-Rodriguez et al., 2020) to extract knowledge from text, markup, and structured sources.
Ontologies are often used to structure human knowledge. Ontology requirements outline what the ontology should be able to do, what knowledge it should contain, and who will be using it (Noy et al., 2001). One of the ways to determine the scope of the ontology is to set a list of questions that a KG based on the ontology should be able to answer known as competency questions (Noy et al., 2001). Following this, conceptualisation techniques help identify and structure the knowledge represented in the ontology, implementation techniques involve selecting the appropriate ontology language and tools to build the ontology, and finally, evaluation techniques assess the ontology’s accuracy, usability, and compliance with requirements (Kotis et al., 2020; Simperl and Luczak-Rösch, 2014).
Quality assessment is the evaluation of the KG on a number of dimensions (Hogan et al., 2022). KGs must be evaluated on their syntactic and semantic accuracy, and their timeliness (Fürber and Hepp, 2011; Martinez-Rodriguez et al., 2020; Esteves et al., 2018; Käfer et al., 2013). Coverage is the KG value of completeness and representativeness (assessing the level of bias of the existing knowledge in the KG, whether this is inherent in the data, the schema or the reasoning) (Martinez-Rodriguez et al., 2020; Darari et al., 2018; Janowicz et al., 2018; Kaffee et al., 2017). Coherency refers to consistency - whether the KG is free of contradiction and consistent with the domain it represents - and validity (Bonatti et al., 2011; Thornton et al., 2019). Succinctness refers to the KG’s conciseness and understandability (by humans) (Hogan et al., 2012; Kaffee et al., 2017).
The above categories can be assessed type using validation tools (Fürber and Hepp, 2011) such as: human evaluation; measuring precision and recall against gold standards or other sources; or comparing known statistical distributions with those of the KG (Martinez-Rodriguez et al., 2020; Esteves et al., 2018; Darari et al., 2018). While not complicated, these assessments can involve a considerable amount of effort. A standard practice to ensure quality and to correct errors is to use rules and constraints that the KG needs to follow (Hogan et al., 2022). Domain experts are the most suitable specialists for creating these rules, however this is an expensive, manual process.
Previous studies argued that knowledge engineers, besides technical skills, require good communication, conceptual ability, patience, organisational skills and knowledge in diverse areas (McGraw and Harbison-Briggs, 1989; Mykytyn Jr et al., 1994). In a collaborative ontology engineering scenario, each member can play several roles, depending on the types of contributions and the technology used (Simperl and Luczak-Rösch, 2014). To merge the gap in communication between domain experts and knowledge engineers, ontology engineering research actively involves domain experts in ontology development with the use of technologies such as controlled natural language, semantic wikis, intelligent user interfaces and social commuting (Denaux et al., 2011). However, all this is disrupted by the advent of LLMs. Today, people and AI agents collaborate to engineer a KG, changing how knowledge engineers work.
2.2. Automation in knowledge engineering
Automation has long held promise for KE tasks. Various automatic processes are utilised for KE tasks, including knowledge extraction, refinement, and enrichment (Tamašauskaitė and Groth, 2023).
Knowledge extraction from structured and unstructured sources (Noy et al., 2019) requires manual and automated approaches. The automated approach can include a number of machine learning techniques: named-entity recognition (NER) to identify name entities in a text, which are words like people, organisations, and locations (Wang et al., 2023); entity linking to link associated entities in a KG with mentions of entities in text (Kandpal et al., 2023); and relation extraction to identify the relation between entities in a text (Cabot and Navigli, 2021). However, automated approaches demand costly and time-consuming techniques with human-annotated datasets for the supervised machine-learning training, and human evaluation for the unsupervised pre-trained models. Furthermore, automated knowledge extraction still presents low accuracy, which can lead to low quality KGs (Veseli et al., 2023; Peng et al., 2023).
KG refinement refers to techniques for KG completion and correction (Liu et al., 2023a; Kumar et al., 2020). For KG completion, the aim is to add missing knowledge to the graph, which is an ongoing process. In this case, KG embedding models are used as a technique to represent information from a KG in a way that computers can understand and work with (Dai et al., 2020). This is the process of turning complex things like entities and relationships in the KG into a simpler format, like numerical vectors. KG embeddings are used for link prediction (Bordes et al., 2013), entity resolution (Ji et al., 2016), relation classification (Weston et al., 2013) etc.
In addition, in KG completion, it may be valuable to incorporate cross-lingual entities, as well as different modalities such as images, videos, audio haptics etc. This is essential for machines to be able to capture the real world in terms of language or human expression through means other than symbolic (Chen et al., 2023). Multilingualism is challenging to capture since there are limited non-English language datasets, entity extraction models must be retrained for each language, and translation systems are not always accurate (Peng et al., 2023).
For multimodality, there are many attempts to describe the different types of knowledge. Some knowledge can only be accurately represented in its native modality, for example, musical chords (de Berardinis et al., 2023). Additionally, while humans are good at providing textual input, some tasks require output in other modes, such as haptics (MuseIT, 2023). The extensive availability of images of the web means that visuals are the most often used media in KGs (Chen et al., 2023). However, finding and integrating other multimodal information can be challenging. There are limited examples of multimodal datasets, making it difficult to find data and ensure that the different modalities refer to the same entity (Chen et al., 2023). These issues are compounded in specific verticals. Knowledge extraction for domains such as heritage remains complicated and time-consuming (van Erp et al., 2023; Cao et al., 2022; de Berardinis et al., 2023).
In contrast, for KG correction, the aim is to correct existing entities and relations. This task can use automatic techniques such as logical inference based on the rules between already existing relations to create and correct new entities and relations. Machine learning techniques can also be utilised for this task such as statistical relational learning, building embedding-based link predictors and node classifiers (Tamašauskaitė and Groth, 2023).
KG enrichment can incorporate ontology refinement and alignment to improve a given KG. In refinement, the aim is to improve text inconsistencies and add meta information like descriptions to entities (Liu et al., 2023a). In addition, for alignment, LLMs have been used to identify and match entities between KGs, for example, using lexical matching (Zhang et al., 2023). In both cases, developing sufficient and reliable tools can be challenging. Traditional methods and techniques require revision and modification to serve knowledge engineers in the production of trustworthy and beneficial KGs.
Opportunities for automation have grown considerably with the advent of generative AI. Pre-trained LLMs like BERT and GPT model families use transformer architecture to achieve state-of-the-art performance for numerous natural language tasks (Min et al., 2023). Two very common processes of LLMs are fine-tuning and prompting (Hu et al., 2023). For fine-tuning, a pre-trained model is further trained on a specific task or dataset to improve its knowledge of a particular task or domain. For prompting, a natural language text is given as input to enable the model to perform specific tasks (Beurer-Kellner et al., 2023). Prompting in a sufficient way to improve models’ outputs requires a number of techniques and rules, introducing the field of prompt engineering (Velásquez-Henao et al., 2023). For prompting, the input to LLM instructions are classified as zero-shot and few-shot prompts, where shot means an example (Velásquez-Henao et al., 2023). Zero-shot prompts are used in situations where it is not necessary to train the LLM, i.e. no specific examples are provided for the task at hand. This relies on the LLM’s ability to follow instructions and generate the desired output. In contrast, in few-shot prompts, the LLM is guided on a task by providing a small set of one or more examples (Velásquez-Henao et al., 2023).
The advent of LLMs revealed two perspectives for KE; KGs can be used to enhance LLMs, and LLMs capabilities can be used to assist on KE tasks (Allen et al., 2023). This study considers the second task.
2.3. Responsible AI
Introducing the use of LLMs to KE requires a close understanding of how to use such AI responsibly. Responsible AI has many underlying principles and themes that ensure the safe deployment of AI systems (Mikalef et al., 2022). Safety in AI is a principle of AI ethics, one that focuses on preventing harm, specifically harm towards human well-being (Kazim and Koshiyama, 2021). AI safety can be used to describe harms in multiple ways. For example, there is the existential safety risk of the fast development of AI without regulation (Bucknall and Dori-Hacohen, 2022), and there is the safe development of current AI systems where AI systems deployed currently hold safety risks and may undermine our trust in them (Shneiderman, 2020). In this study we are interested in the latter.
Safety in AI intersects and is reliant on a number of AI ethical themes (Kazim and Koshiyama, 2021; Khan et al., 2022):
  • Transparency/explainability - being transparent about why a certain decision is made and the capability to explain this appropriately for specific stakeholders.
  • Privacy - ensuring data privacy through the data processing and use lifecycle of AI systems
  • Accountability - ensuring the proper function of AI and compliance with regulation frameworks.
  • Fairness/bias - acting to reduce the likelihood that an algorithm or data systematically produces bias results due to assumptions that are made (Ntoutsi et al., 2020) leading to to unfair outcomes of AI systems that discriminate against certain demographic sub-groups.
LLMs are a black-box models, it is not always possible to understand how outputs come about thus transparency can be very difficult and it is hard to explain their results. This is further complicated by a lack of transparency over data training sets, which also increases the risk of the inclusion of personal data. As generative AI can reproduce data that is present in the training data, it is hard to ensure privacy as it cannot be guaranteed that this data does not appear in outputs. LLMs have been shown to be biased, which can lead to unfair outcomes. Lack of clarity on data and training processes also means accountability is difficult to establish.
Transparency, explainability, and bias are all elements that need considering for any automated KG process; however, they may be exacerbated with the use of LLMs. Previous studies on the use of LLMs show that they lack sufficient accuracy and credibility to be used without human supervision, particularly regarding the lack of provenance and propensity to hallucinate (Ji et al., 2023). Provenance and accuracy are particularly important to KG as a trusted knowledge source, as exemplified by Wikidata. These flag the need for transparency and explainability. Due to the complexity of models, transparency and explainability have been previously explored through documentation for the models or the datasets used to train the models. Mitchell et al. (Mitchell et al., 2019) suggested a framework called Model Cards to encourage transparent model reporting. The framework includes a list of metadata like model details, intended use, factors, demographics, performance, evaluation data, training data, supplementary analysis and ethical considerations. Similarly, Pushkarna et al. (Pushkarna et al., 2022) introduced a framework called Data Cards for purposeful documentation of datasets as an attempt to offer clear documentation of datasets to produce better explained models. The framework suggests a structured list of summaries related to datasets’ metadata such as explanations, rationales, provenance, representation, usage, and fairness.
In terms of bias this may occur in KG using either manual or automatic ways for completion (Voit and Paulheim, 2021). There are also wide ranges of bias, including gender, representatives, and selectivity, which may become encoded into KGs via a number of automated processes. For example, an issue found in information extraction (IE) is the biased statistical dependencies between entities and classes, which is known as spurious correlation (Nan et al., 2021; Ghaddar et al., 2021). Models heavily rely on this type of correlation between the output labels and surface features which effects the performances generalisation. This is due to unbalanced data distributions where entities appear to be casually related but are not. Causal inference is a method that has been used to tackle this specific issue by replacing the target entities (Zeng et al., 2020; Wang and Culotta, 2020, 2021). Counterfactual IE (CFIE) (Nan et al., 2021) extends this method by considering structural information which has been shown to be an important consideration for IE tasks as non-local interactions can be captured (Zhang et al., 2018; Jie and Lu, 2019).
NER has been shown to have bias. There is named regularity bias in which the models’ ability to use contextualised information when predicting the type of entity is ambiguous (Ghaddar et al., 2021). A study that specifically looked at the biases that dictionary definitions have on the outcomes of the NER method (Zhang et al., 2021). NER has also been shown to be better at identifying names from specific demographic groups (Mishra et al., 2020; Paparidis and Kotis, 2021). Demographic biases in KGs have the potential to be caused by neural network-based relation extractions and NER systems excluding protective variables when extracting and processing knowledge (Paparidis and Kotis, 2021). A method where this bias has been addressed is by including information from crowdsourced contributors such as background and information seeking process used to research a certain conclusion, tracking potential biases in collected data (Demartini, 2019).
Semantic role labelling (SRL) is a technique which returns predicate-argument structures for input sentences. This technique has been shown to have demographic bias, such as gender and race bias (Zhao et al., 2017). This technique has also been shown to have commonsense bias where this technique is unable to take into account basic knowledge considered to be commonsense (Lent and Søgaard, 2021). The fact that LLMs are known for generating stereotypes or discriminating data is that they use automated KE deserves particular attention (Nadeem et al., 2021). Further, removal of bias is a difficult task as humans are prone to bias of many types (Keidar et al., 2021; Kraft and Usbeck, 2022).
3. Methodology

### v3

@@ 1 Introduction
    What factors may affect knowledge engineers’ trust and uptake of generative AI technology?
We contribute one of the first user studies on the interaction of knowledge engineers with generative AI. Additionally, we identify key areas of strength and challenge for KE in the era of generative AI. Finally, we propose methods to improve transparency and explainability, including Data Cards for KGs and Model Cards for KG embedding models. It is worth noting that this study was conducted in 2023, and to this day, there is a rapid evolution in the capabilities of generative AI. However, most of the issues raised by participants in the hackathon have not been satisfactorily resolved in the intervening period, and persist as issues that need addressing. Our findings give direction to designers of KE AI copilots to create responsible applications, guide KE researchers and practitioners in developing new methodologies, and advise practitioners wishing to use advanced AI technologies responsibly.
2 Background and Related Works
2.1 Knowledge Engineering
Knowledge engineering, the branch of AI concerned with building and managing knowledge-based systems [150, 164], has changed dramatically with the latest innovations in machine learning (ML), NLP, and computer vision. The process of constructing a KG usually involves acquiring knowledge, processing it, and deploying the knowledge graph [44, 166]. And yet, as the most recent advances in NLP (especially LLMs) and generative AI demonstrate, questions of how to capture and encode domain knowledge in computational representation remain as challenging as ever [147].
Figure 1 shows that the current KG lifecycle consists of four stages. This constitutes a mix of automated and manual capabilities and contributions from several stakeholder groups: knowledge engineers, ML specialists, subject domain experts, online volunteers, and crowdsourcing services, as well as developers of applications using KGs. As the figure suggests, KGs are interacting with AI capabilities in complex ways. Human-in-the-loop tasks in the KG lifecycle increasingly use ML models with varying levels of interpretability.
Figure 1: The KG lifecycle today, illustrating the stages from requirement collection to deployment. Each phase — requirement elicitation, KG construction, maintenance, and deployment — involves collaboration across roles like domain experts, knowledge engineers, and ML engineers to build, enrich, and deploy KGs for intelligent applications.
On the left side of the figure, at stage A, which is an entry point and essential step of the KG lifecycle, knowledge engineers (e.g., ontologists) and KG stakeholders (e.g., domain experts) first determine the scope of work and the success criteria [79]. With the support of description logics, knowledge formalised within a well-structured and expressive TBox that contains terminological axioms can define the relevant vocabulary for a specific application domain. This process, known as ontology engineering (OE), is primarily manual or semi-automatic [12]. Ontology requirement engineering, which serves as the starting point for OE, typically involves writing user stories and deriving competency questions (CQs) from these stories [78].
KG construction generally involves extracting knowledge from multiple heterogeneous data sources, ranging from structured and semi-structured to unstructured data, and integrating it into KGs [166]. At stage B, knowledge engineers and other specialists may reuse standard or novel ontologies to build KGs through data lifting and knowledge extraction. Data from multiple sources is transformed into KGs using ML techniques for tasks such as named entity recognition [181], relation extraction [96] and entity reconciliation [152]. Core KE tasks at this stage include knowledge acquisition and representation. Technologies and end-user tools supporting knowledge acquisition have advanced significantly to meet the scale requirements of modern KGs and lead a growing trend toward adoption of end-to-end approaches [149, 182], which we discuss further in section 2.2. However, the most effective methods for knowledge representation still rely on human oversight at various levels [156, 158]. Increasingly, human input is focused on augmenting or validating algorithmic suggestions [166].
KGs can also be created on a larger scale through human collaboration, utilizing crowdsourcing and collaborative-editing platforms [68]. Crowd workers and editors have important roles throughout the KG lifecycle, especially in knowledge representation and updates, where annotation tasks such as quizzes and voting are often designed to harness their background knowledge [2, 143, 84]. Although KGs developed using these methods may face quality issues, including errors [133, 154], disagreements [85], and bias [68], similar transparency challenges exist in crowdsourcing for supervised ML. These challenges stem from the transparency limitations of the digital services commonly used for such tasks, such as Prolific and Amazon’s Mechanical Turk, which are black-box, proprietary platforms, making results difficult to replicate or reproduce [139]. Furthermore, educating crowd workers to perform annotation tasks effectively is a non-trivial task [143].
On the right-hand side of the figure, KG maintenance (stage C) is prompted by source updates from stage B, and requirements, audits, and assessments from stage D. To further increase their completeness, correctness, and utility, KGs are refined by completion tasks such as link prediction [145] and error detection and correction tasks, etc [28]. Quality assessment of KGs is essential following construction and enrichment stages to ensure they are fit for specific purposes. Key dimensions for evaluation include accuracy, coverage, consistency, completeness and understandability [68]. Accuracy ensures that data represents real-world entities [188], while completeness and coverage assess how well the KG represents all domain-relevant information [31]. Validity assessment is supported by constraints expressed in shape languages, including SHACL and ShEx [51]. By measuring whether the graph conforms to the constraints, knowledge engineers can prevent issues where essential parts of the graph are missing or contradictory. Additional dimensions like coherency, conciseness, and understandability support effective usage by minimizing redundancy and enhancing interpretability, all contributing to the graph’s overall utility for applications [188].
At stage D, KGs support a variety of use cases. They are utilised for querying and reasoning in applications such as search [177], question answering [27, 62], and retrieval-augmented generation [90, 50]. Information can be derived from KGs through deductive (e.g., logical rules) or inductive methods (e.g., continuous graph embeddings) [68]. To ensure trustworthiness and legal compliance, both approaches must be transparent and accountable to users [146, 16].
The roles and required skills in ontology and knowledge engineering have evolved significantly. Beyond technical expertise, knowledge engineers are expected to possess strong communication skills, conceptual thinking, and organizational abilities [103, 112]. In a collaborative ontology engineering, each team member can play multiple roles, depending on the types of contributions and the technology used [157]. To actively involve domain experts in ontology development, technologies such as controlled natural language, semantic wikis, intelligent user interfaces and social computing are proposed [33]. However, the rise of generative AI is reshaping these practices. Collaboration between humans and AI agents is redefining ontology and knowledge engineering, streamlining workflows, and altering the responsibilities of knowledge engineers [8].
2.2 Automation in Knowledge Engineering
Automated methods have long played critical roles in KE, with NLP supporting tasks ranging from knowledge acquisition to completion. The emergence of LLMs has further propelled automated methods into the mainstream, leading to substantial advancements [185]. However, this progress also introduces new challenges in human-AI collaboration and quality assurance.
Once the data source is determined, the construction of KGs typically begins with information extraction, which involves identifying key terms — entities and relations. The task of identifying entities is referred to as named entity recognition (NER). A named entity is a word or phrase that is consistently standard for a real-world thing, e.g., an organization, person, location, etc. The task can be either formulated as a sequence labelling or a span classification task. Current automated NER methods are deep learning-based and consist of input representation at different levels (character, word, hybrid), context encoders that adopt LSTM and Transformers, and tag decoders using models such as conditional random field (CRF) [91].
After determining entities, the next step is to extract relations that link two entities, which is called relation extraction (RE). The task can also be formulated as a span classification task. Given the corpus and a pair of entity strings, the task aims to extract a relation between the entities. Depending on the context given, RE models can be performed at sentence-level, dialogue-level, and document-level. Distant supervision approaches are dominant in current RE tasks, which combine the advantages of prior semi-supervised and unsupervised approaches and leverage a vast amount of semi-structured data stored in a KG [107, 159].
The knowledge extracted in previous steps may contain duplicates, requiring further cleaning and integration. The next step, entity linking (also known as entity alignment, resolution, or matching), is a crucial step of data and knowledge integration. Given two data sources $A$ and $B$, each containing entities with a shared set of attributes, entity linking aims to identify pairs $(e_{1},e_{2})$, where $e_{1}\in A$ and $e_{2}\in B$, such that both refer to the same real-world object. Due to the potential quadratic complexity of matching ($|A|\times|B|$), a pre-processing step known as blocking [125] is often used to reduce the complexity. According to Shen et al. [153], an entity linking system typically consists of three modules: candidate entity generation, candidate entity ranking, and unlinkable mention prediction. Furthermore, PLMs have demonstrated state-of-the-art performance on large-scale, real-world entity linking tasks [92, 162].
The third category is knowledge completion. Most KGs are inherently incomplete, necessitating techniques such as link prediction (also referred to as KG completion or reasoning) to infer missing information. Specifically, given a query $q$ with either missing tail entities $(h,r,?)$ or missing relations $(h,?,t)$, the task is to complete the query by predicting the missing element. Link prediction models are broadly categorised into two main types: rule-based models, for instance, AnyBURL [104], and embedding-based models. Embedding-based models, which represent graph elements in continuous vector spaces, are further divided into conventional approaches (e.g, TransE [20], RESCAL [117], NTN [160], ConvE [35]) and graph neural network (GNN)-based approaches (e.g., R-GCN [148], KBGAT [116][145, 187].
The ability of LLMs to perform knowledge-intensive tasks has been extensively studied, including KG construction and completion [198, 93, 129]. One of the foundational studies, LAMA, examined KG completion by probing language models to extract facts through cloze-style (missing word completion) prompts [132]. Follow-up work after LAMA has shown improvements in performance [140, 198]. TKGCon leverages LLMs to generate high-quality KGs from theme-specific corpora [37]. Veseli et al. [172] conducted a systematic analysis of the capabilities of language models for automated KG completion. Their findings suggest that LLMs can predict facts with high precision for certain relations in Wikidata, though this is not yet universally applicable. Concurrently, prompt engineering techniques have emerged as a focal point in efforts to elicit knowledge from language models effectivelyă[6], further highlighting the evolving role of LLMs in KG-related tasks.
2.3 Responsible AI
Since AI systems are frequently applied in high-stakes contexts such as the medical domain, finance, government, and education, trust concerns have been raised about the potential harm and risks of opaque AI systems to users and stakeholders, especially in deep learning models that are notoriously difficult to interpret [3]. Integrating generative AI into the KG lifecycle may exacerbate existing issues of transparency, explainability, and bias throughout the development and use chain [86]. This underscores the need for a thorough understanding of how to use such technology responsibly. Within the scope of this study, the principles of responsible AI most relevant to KE practitioners include transparency, explainability, fairness, and privacy.
Transparency refers to the degree to which information about an AI system and its outputs is accessible to individuals interacting with the system — regardless of whether they are aware of doing so. It seeks to answer the question of “what happened” in the system. Effective transparency ensures that appropriate levels of information are available at different stages of the AI lifecycle, tailored to the roles and expertise of the stakeholders engaging with the system [4].
Despite its importance, ensuring transparency in the use of generative AI presents significant challenges. Studies indicate that LLMs lack sufficient accuracy and reliability to operate without human oversight, particularly regarding the lack of provenance and propensity to hallucinate [73]. The scale, probabilistic, and inherent black-box nature of LLMs obscure their internal logic, making their behaviour unpredictable and limitations difficult for users to anticipate [61, 196]. Additionally, the opacity regarding which data some of the LLMs are trained on not only reduces explainability but means accountability is difficult to establish and increases the risk of the inclusion of personal data and even the reproduction of such data in outputsă[115]. Given that the training data is obtained from large-scale web crawls [168, 169], LLMs often reproduce and amplify the unfairness and biases inherent in these sources [110, 199].
To address these challenges, several methods have been proposed that emphasise comprehensive reporting of model information, including model details, metrics, evaluation and training data, etc. Mitchell et al. proposed Model Cards, a framework for transparent model reporting that includes metadata on model details, intended use, factors, demographics, performance, evaluation data, training data, supplementary analysis and ethical considerations [109]. It can be tailored depending on the context and the needs of stakeholders, with additional sections such as interpretability methods . Similarly, Pushkarna et al. introduced Data Cards for the transparent and human-centred documentation of ML datasets [138]. A Data Card is a structured collection of summaries of dataset metadata, such as attributes, intended use cases, provenance, along with explanations and rationales. With their high degree of interpretive flexibility, Data Cards can act as boundary objects [161], enabling stakeholders to collaborate by analysing data using shared dimensions and a common vocabulary. Specifically for NLP systems, [15] suggested Data Statements to provide context about datasets, specifically to make NLP users of the dataset aware of issues related to exclusion and biases in the data, or that might emerge with the use of the data.
Achieving transparency in ML models can also be achieved through explainability. eXplainable AI (XAI) encompasses multiple notions, including interpretability, responsibility, and accountability [3]. According to DARPA’s XAI program [^3] https://www.darpa.mil/program/explainable-artificial-intelligence, XAI systems are designed to explain their rationale, characterize their strengths and weaknesses, and help human users to understand “how” they will behave in the future.
Several algorithms and models have been predominently used for XAI. Bach et al. proposed layer-wise relevance propagation (LRP), which decomposes non-linear classifiers like neural networks into several layers of computation and computes relevance scores at each layer that can be used for image classification [13]. LIME, proposed by Ribeiro et al., learns an interpretable model locally around the prediction [144]. Lundbreg and Lee proposed a unified framework that uses SHAP values to measure the feature importance [102]. Certain explainers target specific types of deep learning models. Ying et al. developed GNNExplainer [186], which identifies subgraphs and relevant node features through counterfactual-based importance measurement, applicable across GNN architectures and graph-based ML tasks. With the rise of attention mechanisms [171], attention values are increasingly used for explanations. However, debates persist on their adequacy, with critiques arguing that attention is not inherently explanatory [71] and that more effective alternatives exist [14].
In the context of this study, Danilevsky et al. surveyed state-of-the-art XAI models in NLP [30]. Tiddi and Schlobach’s systematic literature review [167] focused on the integration of KGs into explainable machine learning, where KGs are used as domain knowledge for explanations. Zhang et al. conducted the first examination of XAI methods for KG construction [192]. Beyond the technical scope, Miller et al explore XAI from a sociotechnical perspective, drawing insights from philosophy, cognitive science, and social psychology [106].
Fairness in AI is a key focus in this study, addressing concerns of equality and equity by tackling issues such as harmful bias and unfair outcomes that discriminate against certain demographic groups [4, 120]. Friedman and Nissenbaum defined bias to refer to computer systems that systematically and unfairly discriminate against certain individuals or groups of individuals in favour of others [46]. Bias exists in various forms, including gender, representatives, and selectivity, and may be encoded into KGs through automated processes. Given that LLMs can generate stereotypical bias, their use in automated KE tasks deserves special attention [113].
For example, bias arises in information extraction (IE), NER, and semantic role labelling (SRL). In IE, spurious correlations between entities and classes can skew results, as models often rely on biased statistical dependencies due to unbalanced data distributions [114, 53]. NER is biased in terms of name regularity and demographics, influenced by biased dictionary definitions or neural network-based extraction methods that exclude protective variables [53, 194, 108, 126]. SRL, which analyses predicate-argument structures for input sentences, also exhibits demographic biases, such as gender and race bias [197], and commonsense biases, failing to account for basic commonsense knowledge [89]. However, mitigating bias is challenging, as humans are naturally prone to various biases [77, 86]. Technical approaches to addressing bias include causal inference, which modifies target entities [189, 175, 176], and counterfactual IE (CFIE) [114], which incorporates structural information to capture non-local interactions critical for IE tasks [195, 74].
Beyond the above mentioned principles, Responsible AI encompasses several other key themes [105]. Safety in AI focuses on preventing harm, which includes various themes such as robustness against adversarial attacks, malicious use, reliability and reproducibility, etc [76]. AI safety can describe harms in multiple ways, such as the existential safety risk of the fast development of AI without regulation [25], and the safety risks of current AI systems, which may undermine trust in their deployment [155]. This study specifically focuses on the latter. Privacy ensures data protection throughout the lifecycle of AI systems, both in terms of securing personal information and safeguarding individuals from unwanted exposure [40].
3 Methodology